OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? A) x = 1, x = 17 B) x = −7, x = 9 C) x = 7 x = 9 D) No solution the solutions I got were -1 and -13....

OpenStudy (anonymous):

C) x = 7 x = 9

OpenStudy (anonymous):

how did you get it?

OpenStudy (anonymous):

\[\left| x-8 \right|= \left\{ (8, \infty) \right\} and \left\{ (-\infty , 8] \right\} so \it will be (x-8) \in \left\{ (8, \infty) \right\} and - ( x-8) \in \left\{ (-\infty , 8] \right\}\]

OpenStudy (anonymous):

what? 0.0